chandra has 3 liters of a 14% solution of sodium hydroxide in a container. What is the amount and concentration of sodium hydroxide solution she must add to this in order to end up with 7 liters of 34% solution?
Answer by mananth(16949)

chandra has 3 liters of a 14% solution of sodium hydroxide in a container. What is the amount and concentration of sodium hydroxide solution she must add to this in order to end up with 7 liters of 34% solution?
3 liters 14%
total is 7 liters
so 4 liters x%
..
Total 7 liters 34%
..
Equation for concentration
3*0.14 +4*0.01x = 7*0.34
0.42+0.04x=2.38
0.04x=2.38-0.42
0.04x= 1.96
x= 49% and 4 liters is to be added